OpenStudy (anonymous):

Please help! Will give medal! Ive been stuck on this question for the past 30 minutes: An equation is shown below: 4(2x - 5) + 15 = 11 Write the steps you will use to solve the equation and explain each step.

Mehek (mehek14):

first use distributive property

OpenStudy (anonymous):

So 4(2x) + 4(5) ?@ Mehek14

Mehek (mehek14):

yes that would be the first step

Mehek (mehek14):

so now you have: 8x - 20 + 15 = 11

Mehek (mehek14):

now you would add -20 + 15

OpenStudy (anonymous):

Then subtract 15 on both sides?

Mehek (mehek14):

no first add -20 + 15

OpenStudy (anonymous):

Step 1: Simplify both sides of the equation. 4(2x−5)+15=11 Simplify: (Hide steps) Let's simplify step-by-step. 4(2x−5)+15=11 Distribute: (4)(2x)+(4)(−5)+15=11 8x+−20+15=11 Combine Like Terms: 8x+−20+15=11 (8x)+(−20+15)=11 8x+−5=11 Simplified: 8x−5=11 Step 2: Add 5 to both sides. 8x−5+5=11+5 8x=16 Step 3: Divide both sides by 8. 8x=16 8 8 = 16 8 x=2 Answer: x=2